In a sample of pistons manufactured in a factory, it is seen that 10% of the pistons
are defective. What is the probability that in a batch of 10 pistons no more than two
pistons, are defective?
1
Expert's answer
2020-06-08T10:28:08-0400

The probability of detecting a defective pistons in one test is "p=0.1" then using Bernoully formula "p=p_{10}(0)+p_{10}(1)+p_{10}(2)" where

"p_{10}(0)=\frac{10!}{0!10!}0.1^00.9^{10}\approx0.35 , p_{10}(1)=\frac{10!}{1!9!}0.1^10.9^9\approx0.39"

"p_{10}(2)=\frac{10!}{2!8!}0.1^20.9^8\approx0.19" then "p=0.93"